The hydronium ion concentration of vinegar is approximately 4×10^-3M. What are the corresponding values of pOH and pH?

1
Expert's answer
2022-01-10T13:58:26-0500

pH= - log(H+)pH=−log(H+)


pH= - log ( 4× 10^-3)


pH= 2.4


pOH= 14- 2.4


pOH= 11.6